OpenStudy (anonymous):

Write the equation of a line parallel to y = 3x - 11 and containing the point (2,-3) in y=mx+b format.

OpenStudy (noseboy908):

are you familiar with point-slope form?

OpenStudy (noseboy908):

Also, if a line is parallel to another line (let's pretend with a slope of 1), it has the same slope as that line.

OpenStudy (noseboy908):

therefore both lines have a pretend slope of 1.

OpenStudy (anonymous):

I'm not sure, can you show me how to start the problem please?

OpenStudy (noseboy908):

sure, but do you know point-slope form?

OpenStudy (anonymous):

Yes, y-y1 = m(x-x1)

OpenStudy (noseboy908):

correct. so what's the slope of the original eqn?

OpenStudy (anonymous):

3 ?

OpenStudy (noseboy908):

yes. now plug the new point into point-slope form with the slope from the original eqn, 3.

OpenStudy (anonymous):

So, would it be y-y1=3(x-x1) ??

OpenStudy (noseboy908):

yes. now put in the new ordered pair.

OpenStudy (anonymous):

do i substitue x for 2 and x1 for -3 ??

OpenStudy (noseboy908):

you keep y and x as is. you change x1 and y1 to be the numbers from the new ordered pair: y-(-3)=2(x-(2))

OpenStudy (anonymous):

y+3=3x-6 , is that correct, then subtract 3 from both sides?

OpenStudy (noseboy908):

yep. then you have your answer in m-b form.

OpenStudy (anonymous):

Thank you soo much!:)

OpenStudy (noseboy908):

You're welcome. Enjoy the weekend! Peace out.